OpenStudy (anonymous):

I need some help please. I need to find the intercepts for x and y. can someone explain the steps to me? 8x-7y=168

OpenStudy (anonymous):

set x=0 to get the y intercept, and y=0 to get the x intercept

OpenStudy (anonymous):

so when x=0, y=168/-7 ....... now do the same for y=0, x=168/8

OpenStudy (anonymous):

ok so the ordered pair will be (168,-7) and (168,8)?

OpenStudy (anonymous):

no, your order pair would be (when x=0) (0, -24) (THIS IS WHERE IT CROSSES THE Y-AXIS) (when y=0) (21, 0) (WHERE IT CROSSES THE X AXIS)

OpenStudy (anonymous):

ok so how did you get that?

OpenStudy (anonymous):

ok nevermind i figured it out. thanks